OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of 16 times y cubed − the square root of y cubed?

OpenStudy (anonymous):

√16y^3 - √y^3 = 4√y^3 - √y^3 = 3√y^3 = 3y√y

OpenStudy (anonymous):

I basically first brought 16 out since it's a perfect square, then thought of √y^3 as a variable and combines like terms. Then I brought out y^2 from y^3.